### v1

# Bridging Online Behavior and Clinical Insight: A Longitudinal LLM-based Study of Suicidality on YouTube Reveals Novel Digital Markers

Abstract
Suicide remains a leading cause of death in Western countries, underscoring the need for new research approaches. As social media becomes central to daily life, digital footprints offer valuable insight into suicidal behavior.
Focusing on individuals who attempted suicide while uploading videos to their channels, we investigate: How do suicidal behaviors manifest on YouTube, and how do they differ from expert knowledge? We applied complementary approaches: computational bottom-up, hybrid, and expert-driven top-down, on a novel longitudinal dataset of 181 YouTube channels from individuals with life-threatening attempts, alongside 134 control channels. [^2] Code and data will be available on demand, due to data sensitivity.
[…]
By integrating these approaches, we offer a nuanced understanding of suicidality, bridging digital behavior and clinical insights.
1.   Introduction
Suicide remains one of the most significant public health challenges, particularly among young people in western countries, where it consistently ranks as one of the leading causes of death [34, 17]. Traditionally, research on suicidality has focused on factors such as mental health issues (e.g., depression), social isolation, and family dynamics [13, 39, 14]. These factors were predominantly predefined by experts and were typically examined by relying on patients’ self-reports. However, as social media becomes increasingly prevalent and suicide rates continue to grow, established risk factors and their manifestation are changing, requiring further investigation into suicidal behavior in digital spaces [9, 40].
Social media have become integral to communication and social development, especially among adolescents [35]. Many young people turn to these platforms to share their experiences and seek support [16, 24], potentially providing a digital footprint, documenting the nuances of everyday life, including moments of joy and distress.
The intersection of suicide, social media, and Artificial Intelligence (AI) has gained traction in recent years, with much research utilizing natural language processing (NLP) to analyze large volumes of social media data [37, 25]. However, a common limitation is a reliance on unvalidated suicide risk assessments, such as identifying risk based on messages like "I want to kill myself" or user participation in suicide-related sub-reddits, which introduces biases in both data collection and reliability. Additionally, many studies lack model transparency and have inadequate control groups. Perhaps most importantly, they often fail to incorporate longitudinal analyses that are essential for capturing temporal shifts in relation to suicidal behavior.
Although social media platforms (e.g., Facebook or Reddit) are commonly used to study suicidality, video-sharing platforms like YouTube remain underexplored. As one of the most widely used multimedia-sharing platforms, YouTube enables users to communicate visually and verbally, providing a more direct and unmediated expression. Moreover, YouTube plays a significant role in both shaping and reflecting societal trends related to mental health and suicidality [2, 21]. Still, a search on PubMed revealed only 28 articles containing the words "suicide" and "YouTube" in their title or abstract, with most focusing on the impact on viewers rather than on the creators [11, 22].
Figure 1: An illustration of our research framework, showing the methodological approaches used in this study. Each circle represents a different approach, with the specific methods listed inside. The computational approach includes LLM-based topic modeling to identify behavioral patterns on YouTube (i. §3.1), analyzed through a psychological lens in the discussion section (§4). The domain-expert approach involves a clinical assessment of suicide attempt narratives of predefined psychological factors (ii. §3.3). In the hybrid approach, a domain expert reviewed all the LLM-derived topics and flagged those that may be considered as suicide risk factors, based on current clinical knowledge (iii. §3.4).
This study addresses existing research gaps by bridging the understanding of suicidality in offline and online contexts, considering both behavioral indicators and self-reported experiences, and by incorporating longitudinal analysis of the development of behaviors and feelings before and after suicide attempts. By applying three complementary approaches, as described in Figure 1 and leveraging the strengths of both computational and domain-expert frameworks (Figure 2), we aim to explore our primary research question: How do suicidal behaviors manifest on YouTube, and how are they similar or different from the clinical literature?
Figure 2: Conceptual comparison between the computational approach and the domain-expert approach as applied in this article.
To allow this investigation, we first constructed a novel dataset of YouTube channels belonging to individuals who had attempted suicide (Figure 3), with cases verified as life-threatening by a clinical psychologist (Figure 4). We focused on individuals who attempted suicide after establishing their YouTube channel, referred as Attempted (During). Alongside the Attempted (During) group, which we defined as the treatment group, we gathered three control groups representing varying levels of suicide risk.
Figure 3: The research groups are categorized into treatment - Attempted (During) - and three control groups - Attempted (Before), Control (Matches) and Control (Major Life Event). Both Attempted groups refer to those who attempted suicide during or before their channel upload period, while the other groups refer to control-based matches from the general population, or to those who experienced a major life event during their upload period (and didn’t attempt suicide as far as we know). The sample size for each group is indicated below their respective categories.
In the bottom-up approach, we employed LLM-based topic modeling and analyzed the topic distributions of each video to identify behavioral patterns on YouTube. By examining differences in pre-suicide attempt average topic probability in the treatment group compared to the control groups, we identified potential indicators of suicidal behavior. Then we performed a longitudinal analysis on the identified topics, revealing temporal trends before and after suicide attempts. Through this analysis, we answer the first part of the research question - how suicidal behaviors manifest on YouTube, and partially address the second part, how are they similar or different from the clinical literature, by examining our findings in light of previous suicide research (see Discussion §4).
In the top-down approach, we conducted a clinical psychological assessment of suicide attempt narrative videos to extract relevant predefined factors, including risk and resilience factors. We then statistically compared these factors between YouTubers with a general risk (i.e., attempted suicide in the past) and those at current risk (i.e., attempted during their upload period). This analysis enabled a traditional examination of self-reported suicide narratives.
Lastly, in the hybrid approach, all LLM-derived topics were reviewed by a clinical psychologist to manually identify suicide-related topics, and then subjected to a similar longitudinal analysis as the computational approach. Consequently, this provided a clinical perspective on how the current state of our knowledge about suicide, learned primarily from research in the real, offline world, may appear on YouTube. This approach also expanded the top-down method beyond suicide narrative videos to encompass the full content of YouTube channels.
This methodological framework enables the identification of emerging online behavioral trends and the examination of established psychological factors in a new online context, both of which offer novel insights into suicidality on YouTube (see results summary in Table 2).
Our bottom-up, fully automated computational analysis identified five topics as potential pre-attempt indicators, two of which also exhibited significant temporal shifts around the attempt (§3.1): "Mental Health Struggles" and "YouTube Engagement" (see topics’ representations: Figure 7). "Mental Health Struggles" includes discussions about suicide, self-harm, and depression, aligned with established risk factors. “YouTube Engagement”, which captures typical platform behavior such as audience interaction and content promotion, may serve as an indicator of “normal” YouTube activity, reflecting routine channel maintenance [49, 10, 3]. Notably, this topic appeared at lower levels in the treatment group prior to the suicide attempt—possibly signaling abnormal behavior and, therefore, a potential risk indicator. Both topics exhibited significantly higher prevalence following the suicide attempt, which we discuss at §4.
Additional three topics emerged as pre-attempt indicators and were not associated with a temporal pattern: (1) "LGBTQ+ Identity and Acceptance", (2) "Rape and Sexual Assault", and (3) "Fashion and Style", detailed in §C.3 and discussed in §4. The first two align with established suicide literature, reflecting the heightened distress faced by LGBTQ+ individuals due to discrimination and rejection [46], and by survivors of sexual violence due to trauma-related mental health challenges [12]. Both topics were more prevalent in the attempted groups compared to the other controls. In contrast, Fashion and Style, a less familiar factor in suicide research, showed significantly lower prevalence in the attempted groups, potentially reflecting diminished engagement with self-expression or societal norms.
Moreover, the top-down domain-expert approach showed strong alignment in risk and resilience factors between YouTubers who attempted suicide before and during their upload period, reducing suicide-related tendencies confounders in the comparison between these groups (e.g., depression, family support). The only significant difference was the reason for sharing their attempt story: those who attempted suicide before emphasized their Desire to Help Others, while those who attempted during framed it as part of their Personal Recovery Process (Table 4).
Finally, in the hybrid approach, although 19 topics were flagged as suicide-related by a clinical expert, three were already detected in the bottom-up approach, and only one, “Anxiety and Panic Attacks”, showed a significant group effect, but it did not exhibit a significant temporal change.
The discovery of significant indicators through the computational analysis, which were not identified in either the top-down or the hybrid analysis, underscores the value of bottom-up methods in uncovering hidden indicators. The absence of some expected topics, along with the influence of platform-specific features, such as the YouTube-specific indicators observed in our study, highlights the need for online-focused research. Furthermore, our findings emphasize the importance of distinguishing between current and past suicide attempts, as different patterns may emerge across these groups, reflected in both the computational and expert-driven analyses.
The key contributions of this study are: (1) novel insights into the manifestation of suicidal behavior in online contexts, including reduced engagement-seeking content as a strong behavioral marker; (2) integration of clinical and computational perspectives, revealing both aligned and distinct indicators; (3) uncovering psychological similarity between individuals who attempted suicide before and during their YouTube activity, differing mainly in their motivation for disclosure; (4) creation of a first-of-its-kind publicly available longitudinal dataset of 181 suicide-attempt channels and 134 controls, enriched with transcripts, metadata, demographics, and psychological assessments; (5) development of an AI-driven pipeline for analyzing longitudinal YouTube content; and (6) a novel matching framework for assembling high-quality YouTube control cohorts. We hope this work will inspire future efforts to integrate digital behavioral analysis with clinical insight in the study of suicidality.
2.   Methods
This section outlines the methods used in this study, beginning with data collection (Figure 4) and followed by a description of the procedures (Figure 6). The data collection2.1.1) involved constructing a textual dataset comprising four types of YouTube channels creators (Figure 3): belonging to creators who attempted suicide (before or during their YouTube activity), those who experienced a major life event, and matched control users from the general population. We searched YouTube for relevant channels, transcribed the videos and collected channel information, annotated demographic and psychological assessments, and matched them to control channels using a multi-step approach. The procedures2.2) included applying an LLM-based topic modeling to the channels’ transcripts and implementing statistical procedures to address our research objective: understanding how suicidal behavior manifests on YouTube.
2.1 Data
2.1.1 Data collection
YouTube served as our primary data source. We identified users who uploaded a video sharing either their own suicide attempt story or a major life event using the YouTube API. [^4] API Documentation: developers.google.com/youtube/v3/docs These stories were typically not the main theme of the channels, but were often shared to raise awareness of suicide or loss. Additionally, we included a control group that matched the treatment users to facilitate a fair comparison with the general population. The data collection procedure is outlined in Figure 4 and described in more detail below, with full documentation—including inclusion and exclusion criteria—provided in the Supplementary Materials (Table A.1).
Figure 4: The data collection pipeline includes searching for relevant YouTubers (e.g., YouTubers who attempted suicide), a two-step human evaluation, involving non-expert assessment (personal channel validation and extracting demographic and event-related information), and psychological assessment (attempt intent validation and suicide-related factor extraction); both steps validating and then analyzing the narrative stories describing a suicide attempt or a major life event. The final steps involved control-based matching algorithm, which involved finding "similar" YouTube channels to the treatment group (Figure 5). The last step included performing data processing (including filtering uploads and automatic transcription) on the 4 research groups.
First, using the YouTube API, we searched channels of people who attempted suicide (e.g., ’my suicide attempt’) and people who experienced a major life event, mostly involving the loss of a close relative. These disclosure videos are named narrative videos.
We then conducted a human evaluation involving two levels of expertise. First, non-expert channel validation to ensure that only personal channels were included, while excluding educational, news, and family channels. Following this, we extracted user and event information from the attempt and major life event stories (e.g., attempt date). Second, a psychological assessment of the attempt narrative videos of YouTubers who reported a suicide attempt for two purposes: (1) validating the intent of the attempt, and (2) comparing Attempted (During) and Attempted (Before) to suicide-related tendencies to estimate potential confounding factors (e.g., mental illness, risk groups) and to perform a suicide-related factor analysis.
Next, we applied a multi-step matching process (Figure 5), which allowed us to first search for control cases from the general YouTube population and later match them to users who attempted suicide, based on demographic and channel characteristics.
Finally, we employed data processing, which included collecting the users’ channel videos and metadata (e.g., likes, description) and performing automatic transcription using Whisper [41], the current speech-to-text state-of-the-art model.
Figure 5: The matching algorithm, comprising of four steps: (1) creating queries based on the treatment group, using the YouTube API (§A.4.1); (2) automatic and manual filtering resulting in relevant personal channelsA.4.2); (3) propensity score-based matching based on demographic and channel data, using (a) logistic regression and (b) KNN matching of treatment to control users (§A.4.3); and (4) LLM-based refined matching, using GPT-4 to score the potential matching, retaining good matches (scores > 3) (§A.4.4)
2.1.2 Study population
We categorized channels into two primary groups: treatment and control (Figure 3). The treatment consisted of YouTubers who attempted suicide during their upload period, while the first control group included YouTubers who attempted suicide before their upload period. These groups were labeled as Attempted (During) and Attempted (Before). The additional control groups comprised YouTubers who experienced a major life event during their upload period, labeled Control (Major Life Event), and YouTubers matched to the treatment group based on demographic and content characteristics, labeled Control (Matches). The control groups consist of both a potential at-risk population that had encountered major life events or attempted suicide in the past, and a general population resembling the treatment users. See Table A.1A.1) for inclusion and exclusion criteria.
Group # Users Uploads (Total) Uploads
(Mean / Med / SD)
Female
(n, %)
Age (Mean) Minority (n, %)
Attempted (During) 42 6,354 151 / 89 / 169 30 (71%) 24 27 (59%) Attempted (Before) 139 28,053 201 / 83 / 312 91 (65%) 26 78 (56%) Control (Major Life Event) 43 13,470 320 / 155 / 372 29 (69%) 27 15 (35%) Control (Matches) 95 14,860 182 / 95 / 223 79 (81%) 23 52 (54%)
Table 1: Descriptive Statistics of the Research Groups. Summary of demographics and YouTube activity. See §A.6 for upload trends pre- and post-event.
The Final Sample (Table 1): Of the initial 303 YouTubers in the Attempted group, 104 users were excluded due to non-suicidal cases (e.g., clickbait-like videos like "Bust Myths About Suicide"), non-English content, or an insufficient number of videos, and 18 users were reclassified as Control (Major Life Event) due to the suicide attempt of a close relative. The final set of attempted channels consists of 181 channels (66% female, average age = 25) who uploaded 34,407 videos ($M=190$, $Mdn=83$, $SD=286$). This dataset was later divided into two groups: 42 Attempted (During), defined as the treatment, and 139 Attempted (Before), defined as one of the three controls, based on whether the attempt occurred during or before their upload period (Figure 3).
A similar filtering process was applied to the Control (Major Life Event) group. The final Control (Major Life Event) dataset included 43 users (69% female, average age = 27). These users collectively uploaded 13,470 videos ($M=320$, $Mdn=155$, $SD=372$). This subgroup included 25 users who experienced prolonged difficulties (e.g., relative’s cancer), with the remaining 18 users experiencing the acute death of a family relative (e.g., heart attack).
The final Control (Matches) group included 95 channels (81% female, average age = 23) that met the inclusion criteria and were found as most similar to the 42 Attempted (During) group according to the matching algorithm (Figure 5). These 95 users collectively uploaded 14,860 videos ($M=182$, $Mdn=95$, $SD=223$).
2.2 Procedures
This study introduces a novel longitudinal analysis of linguistic content from YouTube channels, focusing on suicide attempts as a reference point for temporal changes. The procedures pipeline is visualized in Figure 6, and outlined below.
Figure 6: The methodological pipeline consists of three approaches: 1. Computational Bottom-Up - Factors were generated using an LLM-based topic modeling algorithm, and topic relevance was evaluated using a bottom-up approach; 2. Domain-Expert Top-Down - Factors were predefined by a clinical expert and assessed solely based on the suicide attempt narratives, and a 3. Hybrid Approach - A set of 166 LLM-generated topics was reduced to 19 based on their relevance to suicide as determined by a clinical expert. The bottom-up and hybrid approaches were also followed by a two-way mixed effect ANOVA examining both temporal and group effects, while the top-down approach was limited to group differences.
LLM-based Topic A topic is defined as a cluster of semantically related text segments gathered by an unsupervised clustering method of sentence embeddings. Each topic is automatically titled by LLM considering top-associated words and representative texts.
For topic modeling details, see §B.1.
First, we applied an LLM-based topic modeling technique (BERTopic [18],
Figure B.1). To capture the language and behavioral patterns typical of the at-risk population, the topic model was trained exclusively on the textual content from the Attempted users’ channels, comprising 757,849 text segments, yielding 166 topics (see Definition 2.2). This model was then used to infer topic distributions across the videos of all research groups (Figure 3). These topic distributions are used in bottom-up and hybrid approaches.
Next, we defined the reference event of each group: a suicide attempt for Attempted (During), a major life event for Control (Major Life Event), and a synthetic event (the midpoint of uploads) for Attempted (Before) and Control (Matches). Topic probabilities were then averaged separately for the pre- and post-event periods. Although the attempt narratives were included in the topic model, they were excluded from subsequent statistical analyses to avoid bias, as they do not represent the typical content of the channels.
The bottom-up and hybrid procedures followed a two-step protocol: topic selection, aimed at enabling more reliable statistical analysis in a high-dimensional setting, and a statistical effect analysis to evaluate group and time effects.
In the hybrid approach, we asked a clinical expert: What topics are known risk factors of suicide? which reduced the topics from 166 to 19. In the bottom-up approach, we asked: What topics may serve as indicators of the period before a suicide attempt?, and used stepwise logistic regression to the pre-event period, coding Attempted (During) as 1 and controls as 0, across three control populations and their combined pool, yielding four binary models and narrowing topics from 166 to five.
Following topic selection, we conducted a two-way mixed-effects ANOVA [4], on the time-aggregated topic probabilities, in which the group type was the between-subjects factor, and time (pre- or post-reference point) was the within-subjects factor. Based on the significant effects identified in the ANOVA, we performed Tukey’s post-hoc tests [1], to elucidate group differences and temporal patterns. See §C.3 for detailed statistical analysis.
As a complementary analysis, we extended our time-aggregated approach by conducting two types of longitudinal topic analyses to explore temporal patterns of significant topics. The first was a within-group analysis, where we explored changes over time within the Attempted (During) group. This analysis used $t=0$, allowing us to provide a standardized baseline from which to compare temporal changes. The second was a between-group analysis, in which we compared the Attempted (During) group to the three control groups at each timestep.
Finally, motivated by traditional clinical assessment practices, we conducted a top-down analysis in the form of a psychological factor analysis, aiming to assess whether the behavioral signals identified in our study align with established clinical knowledge on suicidality. This analysis is based upon assessments conducted by a clinical psychologist on the Attempted groups (§2.1.1) and involved one-hot encoding of the psychological factors, then performing a stepwise logistic regression analysis [38], where the outcome variable was coded as 1 for Attempted (During) and 0 for Attempted (Before).
3.   Results
Based on the research framework illustrated in Figure 1, this chapter presents the results of the three approaches (Table 2): bottom-up computational (§3.1), top-down domain-expert (§3.3), and hybrid (§3.4), complemented by longitudinal topic analysis (§3.2).
[…]
3.1 The bottom-up analysis
Our primary computational analysis examines LLM-generated topic distributions (Figure C.1). Statistical comparison between the treatment Attempted (During) group and the control groups found the two topics - YouTube Engagement and Mental Health Struggles, as both predictive of the pre-attempt period and significant indicators of temporal changes (pre to post-attempt). Representations of these topics are presented in Figure 7.
Figure 7: Word clouds generated from TF-IDF scores, highlighting the importance and frequency of words in the two topics that significantly predicted the pre-suicide attempt period and exhibited notable temporal changes across groups. The size of each word corresponds to its TF-IDF score. Below each topic, a representative transcript segment is presented.
3.1.1 What topics may serve as indicators of the period before a suicide attempt?
As part of the topic selection procedure, we identified topics predictive of the period preceding a suicide attempt with a forward-backward stepwise logistic regression [38], as these factors may act as indicators of current risk for suicide attempt.
The analysis revealed five significant topics (Table 3): (1) YouTube Engagement, (2) Mental Health Struggles, (3) Fashion and Style, (4) Rape and Sexual Assault and (5) LGBTQ+ Identity and Acceptance. Figure C.3 illustrates the word cloud of these topics, and Figure C.4 illustrates their temporal dynamics.
Treatment Group Control Group Topic $\beta$ Coeff. $z$ Score $p$ Value 95% CI Attempted (During) Attempted (Before) Mental Health Struggles -1.49 -3.62 $<0.001$ [-2.29, -0.68] Rape and Sexual Assault 0.68 2.68 0.007 [0.185, 1.186] Control (Major Life Event) YouTube Engagement -1.46 4.23 $<0.001$ [-2.142, -0.787] LGBTQ+ Identity and Acceptance 0.96 2.58 0.010 [0.233, 1.7] Control (Personal) YouTube Engagement -0.808 -3.19 0.001 [-1.304, -0.313] Fashion and Style -0.85 -2.74 0.006 [-1.462, -0.243] Combined Pool YouTube Engagement -0.73 -3.92 <0.001 [-1.098, -0.366]
Table 3: Regression Result: Each model predicts membership in the Attempted (During) group versus either an individual control group or a combination of all three (i.e., Combined Pool), using pre-event average topic probabilities as features. The analysis yielded four binary logistic regression models, each focusing on a different experimental setup.
3.1.2 Temporal effect analysis of pre-attempt indicator topics
Following the topic selection procedure (§3.1.1), which identified five topics, we conducted a two-way mixed ANOVA to examine the effects of group and time, as well as their interaction, across the four research groups. As shown in Table C.1 and illustrated in Figure 8, significant group-by-time interactions were observed for the topics YouTube Engagement and Mental Health Struggles. As illustrated in Figure 8, the Attempted (During) group exhibited a post-event increase in both topics, aligning with the Attempted (Before) group. For YouTube Engagement, both Attempted groups (Before and During) showed lower levels before and after the event compared to the control groups.
Figure 8: YouTube Engagement and Mental Health Struggles. The figures present the average topic values across four groups: Attempted (During), Attempted (Before), Control (Major Life Event), and Control (Personal), comparing Pre- and Post-Event (suicide attempt, major life event, or synthetic condition). The topics YouTube Engagement and Mental Health Struggles were first identified as significant in the bottom-up analysis (§3.1). The word clouds of these topics are presented in Figure 7.
These results were further examined using post-hoc Tukey’s HSD tests. Group-level differences were assessed via pairwise comparisons of overall topic means across groups (Table C.2), identifying specific group pairs with significant differences in topic prevalence. To investigate the group × time interaction effects, additional post-hoc analyses were conducted by comparing the temporal change (i.e., pre- vs. post-event differences) across groups (Table C.3). This analysis found that the Attempted (During) group showed significantly different temporal changes in YouTube Engagement compared to all control groups, and in Mental Health Struggles, only compared to the Attempted (Before) group.
3.2 The longitudinal topic analysis
Figure 9: Longitudinal Comparison of YouTube Engagement (left) and Mental Health Struggles (right) relative to $t=0$ in the Attempted (During) Group. The figures depict temporal trends of these topics within the Attempted (During) group, with the timeline aligned to the suicide attempt. At each time step, comparisons were made to $t=0$ using the paired t-test. Asterisks (*) indicate statistically significant differences. The full statistical procedure is detailed in §B.3.
We conducted a longitudinal analysis focusing on the topics Mental Health Struggles and YouTube Engagement, as they exhibited significant temporal changes following a suicide attempt (§3.1.2).
Figure 9 shows the results of the within-group analysis, which illustrates how the prevalence of YouTube Engagement and Mental Health Struggles significantly increased after the attempt. Additionally, the between-group analysis, depicted in Figure 10, reveals that YouTube Engagement significantly differs between the Attempted (During) group and all control groups in the time-steps right before the attempt. Meanwhile, significant differences in Mental Health Struggles were primarily observed between the Attempted (During) and Attempted (Before) groups. These differences were further validated as relative to suicide attempt, as a control analysis using COVID-19 as a reference event showed no significant differences between-group effects (§C.3.3).
Figure 10: Longitudinal Comparison of YouTube Engagement (left) and Mental Health Struggles (right) Between Attempted (During) and Control Groups. The figures illustrate temporal trends of these topics across the four groups. The timeline is aligned to a unified event-based reference point (suicide attempt, major life event, or synthetic condition). At each time step, the Attempted (During) group was compared to the other groups using the t-test. Asterisks (*) denote statistically significant differences. The full statistical procedure is detailed in §B.3
3.3 The top-down domain-expert analysis
While an AI-driven bottom-up approach uncovers topics without relying on clinical knowledge, traditional clinical assessments provide structured interpretations grounded in psychological literature. To combine these perspectives, we incorporated a domain-expert top-down analysis. In this analysis, a clinical psychologist assessed psychological factors by analyzing only the suicide attempt narrative videos, whereas the bottom-up approach was applied to the entire channel. This method allowed us to construct a psychological profile of the attempted groups, which we later incorporated into a regression analysis to differentiate between the two groups.
We found that the Attempted (During) and Attempted (Before) groups reported similar risk and resilience factors in their suicide narratives (Table 4). Correspondingly, the users in both the Attempted (During) and the Attempted (Before), display a similar distribution of mental health conditions (e.g., 78.6% and 79.9% self-reported depression), resilience factors (40.5% and 37.4% self-reported social/family support), and experiencing feelings of hopelessness/despair (71.4% and 64%) and unbearable mental pain (38.1% and 46.8%). The only significant difference between the groups lies in their attempt sharing motivations: Attempted (During) users mainly discuss Personal recovery (66.7% vs. 38.1%), while Attempted (Before) users mentioned the Desire to help others (77% vs. 40.3%).
[…]
Since the main bottom-up approach was not specific to the suicide attempt narratives, we conducted an additional analysis using a separate LLM-based topic model focused solely on suicide attempt narratives (§C.2). This complementary analysis enabled us to identify thematic patterns within the narratives, and initial exploration shows alignment between LLM-generated topics and predefined suicide-related psychological factors.
Category Sub-Category Attempted (During) (%, n=42) Attempted (Before) (%, n=139)
Risk Groups
Teens 38.1 47.5
Immigrants 33.3 27.3
Men 26.2 26.6
LGBTQ+ 11.9 13.7
Suicide Survivor Family 0 2.9
Mental Illness
Depression 78.6 79.9
Anxiety 31 30.2
Eating Disorder 14.3 5.8
PTSD 2.4 8.6
Substance Abuse 4.8 5.8
Video Motivation
Desire to Help Others 38.1 * 77 *
Personal Recovery Process 66.7 * 40.3 *
Raising Awareness 21.4 20.1
Seeking Help 2.4 1.4
Resilience Factors
Social/Family Support 40.5 37.4
Responsibility for others 14.3 16.5
Treatment 14.3 15.8
Fear of Death 2.4 11.5
Self Disclosure Ability 9.5 7.9
Risk Factors
Hopelessness/Despair 71.4 64
Loneliness 47.6 59
Unbearable Mental Pain 47.6 59
Life Events 45.2 43.9
Comorbidity 19 25.2
Impulsivity and Aggression 26.2 20.1
Self-Harm 21.4 18.7
Bullying 9.5 23
Access to Lethal Means 26.2 17.3
Self Disclosure Difficulties 19 18
Lack of Belonging 19 18
Abuse 16.7 16.5
Cognitive Distortions 4.8 5.8
Suicide Attempt Validation
Suicide Attempt 100 86.3
Suicidal Ideation 54.8 70.5
Non-Suicidal Self-Harm 0 1.4
Table 4: Psychological Factors Analysis conducted by a clinical psychologist through narrative analysis of attempt stories. The assessment comprised six sections: Risk Groups, Mental Illness, Video Motivation, Resilience Factors, Risk Factors, and Suicide Attempt Validation. The table displays the percentage of individuals in each group, with each condition determined by the psychologist based on the content described in the attempt stories. The data indicates "similar confounding" factors across most features, except for the motivations for posting the video, marked with *, and validated as significant. The table includes only factors that could be extracted from a sufficient number of users (>5%). See §A.3 for the assessment protocol and §C.2, which describes an LLM-based topic model trained on the narrative stories.
3.4 The hybrid bottom-up and top-down analysis
The bottom-up approach automatically generated LLM-based topics, which were analyzed computationally and longitudinally to identify behavioral indicators of suicide attempts (§3.1). In contrast, the top-down approach involved manual assessment of pre-defined psychological factors and analysis of suicide attempt narrative videos (§3.3). To integrate both approaches, we asked a clinical psychologist to evaluate all 166 LLM-derived topics (2.2) and flag those associated with suicidality. We then conducted the same statistical effect analysis used in the bottom-up approach on this expert-curated subset (§3.1.2).
3.4.1 What topics are known risk factors for suicide?
The clinical psychologist identified 19 topics as related to suicidality based on topic descriptions, three of which were already identified by the computational approach (§3.1). In this section, we focus on the remaining 16 topics. The full list is available in §D.
3.4.2 Temporal Effect Analysis of Suicide-Related Topics
Figure 11: Anxiety and Panic Attacks. The left figure presents the average topic values across our four groups, comparing Pre- and Post-Event (suicide attempt, major life event, or synthetic condition). The right figure presents the topic’s word cloud.
[…]
As shown in Figure 11, the Attempted (During) and Control (Major Life Event) groups had lower topic prevalence before the event, with some increase observed post-event. However, post-hoc tests (Tables C.5C.6) indicated that the Attempted (Before) had a significantly higher group prevalence, and the observed temporal increase in the Attempted (During) group was not statistically significant.
Interestingly, well-established clinical suicide-related factors, such as Religion, Addiction Recovery, and Bullying (Table LABEL:tab:topic_info), did not yield statistically significant effects in our analysis, despite being flagged as relevant topics by a clinical expert. Additionally, other risk factors, such as Hopelessness or Social Isolation, did not appear in the LLM-derived topics.
4.   Discussion
This study introduces a unique, publicly available longitudinal dataset of 181 YouTube channels from individuals with verified suicide attempts, along with two additional control groups (n = 138), analyzed through a multi-method framework combining bottom-up, top-down, and hybrid approaches. Anchoring the analysis around the suicide attempt, the bottom-up computational approach identified Mental Health Struggles and YouTube Engagement as key behavioral markers, both significantly lower before the attempt compared to control groups and increasing afterward. The top-down psychological assessment of suicide narratives confirmed the suicide attempts as life-threatening, and revealed strong alignment between the narratives of Attempted (Before) and Attempted (During) groups in psychological factors such as depression and loneliness. The only notable difference was motivation to share their suicide attempt, with the latter framing it as part of a Personal Recovery Process. Lastly, the hybrid approach identified 19 out of 166 LLM-derived topics as suicide-related. However, only Anxiety & Panic Attacks was additionally supported by statistical analysis. Together, these complementary perspectives provide a richer, more nuanced understanding of how suicidality manifests on YouTube.
A particularly intriguing finding was the increase in the discussion of Mental Health Struggles after a suicide attempt. This finding is in line with previous studies indicating that mental health struggles may increase after a suicide attempt, since mental health disorders (depression, anxiety, PTSD) are often persistent or worsen post-attempts [29]. Moreover, the period after an attempt is a high-risk window for another attempt [29]. Individuals after an attempt may feel increased feelings of shame and guilt, and their interpersonal relationships, may be challenged [45]. However, this observation might appear counterintuitive based on some traditional psychological expectations, which suggest that discussions about mental health struggles typically increase prior to an attempt, as individuals at risk are more likely to seek help from mental health services [26], and may diminish post-attempt due to a sense of relief or emotional avoidance [6, 20].
Assuming the latter theory holds, possible explanations for the discrepancy between this theory and our findings may relate to the nature of the data we rely on or may be explained by psychological theory. First, consider the characteristics of the study population and platform. As noted by [26], “…alternate approaches to suicide-prevention efforts may be needed for those less likely to be seen in primary care or mental health specialty care, specifically young men.” Our research population differs from those typically studied, focusing on young people and YouTube creators in the context of their online behavior. Moreover, YouTube content may not fully capture the breadth of an individual’s emotional experiences, as video creation depends on both the technical and emotional readiness of the creator. Second, the clinical literature suggests that the increased descussion on mental health could be related to the process of post-traumatic growth [15]. Individuals who survive a suicide attempt may experience psychological growth through sharing their struggles [23]. This hypothesis is further supported by our psychological analysis, which found that the Attempted (During) group was significantly more motivated to share their stories as part of a Personal Recovery Process.
The second main finding, regarding YouTube Engagement (content aimed at driving audience interaction), is an intriguing behavioral marker, as it may reflect normal platform usage patterns rather than being directly indicative of distress. Given that creators often seek interaction to maintain their audience [49, 10, 3], this finding underscores the importance of contextualizing social media behaviors within platform norms. The finding regarding higher YouTube engagement after a suicide attempt may be explained in several ways. Engaging others may be a form of seeking connection. Individuals often feel isolated after a suicide attempt and turn to others in order to understand and validate their struggles and feelings. A social media community can provide empathy and reassurance. Another option is that the engagement is actually a cry for help [27]. This may be a way to signal that they are still struggling and need support. They might not feel comfortable asking for help directly, but hope that someone will notice and intervene.
Additional findings emerged regarding three topics identified as potential pre-attempt indicators in the computational approach: (1) LGBTQ+ Identity and Acceptance, (2) Rape and Sexual Assault, and (3) Fashion and Style. The first two align with existing literature on suicide risk factors and were independently identified as suicide-related by a clinical psychologist in the mixed-method approach. LGBTQ+ individuals are at higher risk for suicide attempts compared to the general population [43, 8]. The chronic stress experienced by LGBTQ+ individuals due to social rejection and discrimination is extremely high [46]. In addition, LGBTQ+ individuals face many barriers to mental health care [32]. Similarly, Rape and Sexual Assault are profoundly traumatic experiences that elevate the risk for attempts [12]. These traumas often lead to mental health disorders, including depression and substance abuse, which are themselves risk factors. Moreover, the victims often suffer from social isolation and stigma, which intensifies feelings of loneliness and despair. In contrast to the first two topics, Fashion and Style, a less familiar factor in suicide research, showed a significantly lower prevalence in the attempted groups. This may reflect reduced engagement with self-expression or social norms. Prior studies link between mood and clothing and suggest that fashion may indicate on positivity [33, 28]. This pattern may also align with Joiner’s Interpersonal Theory, where social withdrawal and loss of belonging contribute to suicidal behavior [19]. The decrease in fashion-related content may indicate broader patterns of social disengagement and loss of interest in activities that typically foster interpersonal connection. However, given that Fashion and Style are not conventionally examined in suicide research, further studies are needed to clarify this relationship.
Despite these promising findings, this study has several limitations. First, the sample is inherently biased toward individuals who are active on YouTube, which may not be representative of the broader population at risk for suicide. Additionally, the reliance on public content means that more private or intimate indications of distress may not be captured (as would be the case in, e.g., clinical records and personal messages). Another limitation is that all clinical annotations were performed by a single expert, which may introduce subjective bias. This limitation stems from the demanding nature of the task, which required reviewing and coding hours of suicide narratives. Future studies should consider incorporating data from other platforms and employing methods to assess more private, less curated expressions of suicidality.
Looking ahead, future research could explore and compare how suicidal behaviors manifest across different social media platforms. A particularly compelling direction would be to examine engagement techniques across platforms and analyze their connection to mental health-related behaviors, as YouTube Engagement is identified in this research as a strong and consistent behavioral indicator. Another promising direction is leveraging YouTube’s multi-modal content, including speech, audio, and visual, to enhance predictive power and capture nuanced behaviors that may have been overlooked in this study.
In summary, our study adds to the growing body of literature on social media and mental health while also advancing the use of YouTube for suicide research. By integrating qualitative and quantitative methods and emphasizing longitudinal analysis considering suicide attempts, we highlight the significance of a temporal approach compared to broader group-based analyses. We hope this work encourages further interdisciplinary research in this critical area.
Appendix A Appendix A - Data Collection
A.1 Inclusion and exclusion criteria
Stage (Scope) Filtering Step Group
Attempted (During) Attempted (Before) Control (Major Life Event) Control (Matches)
Search (Video) Search Queries (OR)
“my suicide attempt”
“my suicide story”
“I am a suicide attempt survivor”
“attempted suicide”
“take my life”
“my relative died”
“losing my relative
“I lost my relative
[…]
father, mother}
Automated channel- specific queries using: • Frequent bigrams • Video categories • Upload periods
Search (Video) Category
Excluded query videos from the following categories:
["Gaming", "Sports", "Comedy", "Music", "News & Politics",
"Pets & Animals", "Science & Technology", "Travel & Events", "Film & Animation"]
Search (Video) Content Excluded query videos that appeared to be: • clickbait (e.g., "attempt… [Mha gacha] [Warning:Suicide]") • news reports (e.g., "Coco Lee dies after suicide attempt") • non-personal stories (e.g., "TONI Ep.104 | Molested and Attempted Suicide")
Search (Channel) Personal Accounts Included personal accounts validated via profile picture, channel link and sample video
Preprocessing (Video) Content Included Only public, English, downloadable videos, longer than 60 seconds.
Preprocessing (Video) File Size Limit Not applied Excluded $>$50MB
Preprocessing (Transcript) Transcripts Excluded transcripts if: (1) $<$3 sentences, (2) $>$50% repeated (hallucinated) content, (3) $>$30% non-English, or (4) $>$2 speakers. Videos which weren’t filtered are considered valid.
Preprocessing (Channel) Activity Threshold Excluded channels with $<$10 valid videos
Assessment (Mixed) Group-Specific Procedure Psychological assessment + Event validation Event validation Matching Procedure
Final Sample Size n=42 n=139 n=43 n=95
Table A.1: Inclusion and exclusion criteria for the four research groups. The leftmost column indicates the pipeline stage and data scope at which each filtering step was applied (e.g., during search, query filtering, or preprocessing). Label Studio was used for manual review.
As part of the data collection, we used the YouTube API ,Google YouTube API Documentation, to search for channels of people who are relevant for our research, i.e., those who attempted suicide - Attempted (During) and Attempted (Before), and those who experienced a major life event - Control (Major Life Event). See Table A.1 for full inclusion and exclusion criteria for the research groups.
A.2 Data preprocessing:
During the data collection §2.1.1, we implemented the following preprocessing steps:
  1. 1.
    Downloading: We downloaded all channel uploads exceeding 60 seconds (no upper limit on video duration), excluding unavailable videos (e.g., private uploads). The downloading process, which spanned several months, utilized the yt-dlp package to obtain videos in .mp3 format and bright-data services for port forwarding services.
  2. 2.
    Whisper Transcription: We used the Whisper (medium) [41], model for audio-to-text conversion with specific parameters to enhance transcription accuracy (no speech threshold=0.3, condition on previous text=True).
  3. 3.
    Data Cleaning: To ensure dataset quality and relevance, we excluded transcripts that match one of the following criteria :
    1. (a)
      Too short - fewer than three sentences.
    2. (b)
      Contain hallucinations - repeated words or sentences comprising more than 50% of the document, often resulting from non-verbal content like background music [50].
    3. (c)
      Non-English Content - contain more than 30% non-English content (detected automatically).
    4. (d)
      Non-Personal - more than 2 speakers (detected automatically).
    The remaining videos were labeled as valid.
  4. 4.
    Transcript Segmentation: To accommodate the token limits of modern embedding models like Sentence-BERT [42] (which has a 512-token threshold), we divided transcripts into chunks of a maximum of 80 tokens, ensuring each chunk contained complete sentences. This segmentation allows for better representation of multiple topics within a single video.
A.3 Human evaluation
As part of our annotation process, we used the Label Studio [47] platform to create two protocols. In both, the annotators were provided with a YouTuber image, Link to the YouTuber’s channel, Attempt/Major Life Event narrative’s video, accompanied with audio, and transcription of the video, and summary (automatically generated by GPT-4 [36]).
The Non-Expert Assessment step covered three broad areas: demographic information (e.g., gender, age, ethnicity), channel-based features (e.g., confirmation of first-person accounts, personal channels), and event-related details (e.g., method of suicide attempt, such as medication, for attempts; or type of major life event, e.g., sudden death of relative). In addition, the date of the suicide attempt and the major life event was extracted as part of the event-related information, and defined as the reference points. The dates were recorded, with some users providing specific dates (e.g., 02-05-2017) and others providing approximate estimates (e.g., beginning of the year, October 2017). Not all groups had a specific event date during their upload period (i.e., Attempted (Before) and Control (Matches)), thus, we also created synthetic reference events for different groups (e.g., middle of the uploading period). These reference events allowed us to control for the significance of the suicide attempt compared to other reference events.
This annotation was performed on Attempted (During), Attempted (Before), and Control (Major Life Event).
The annotators were asked to answer the following questions:
  • Relevance -
    • What type of channel is it? Personal (individual who uploads content about themselves) or General (e.g., family channel, news outlet, educational content)
  • Demographic Information - age, ethnicity, gender.
  • Event information -
    • Does the story talk about a suicide attempt/major life event?
    • Who did the attempt/died?
    • What is the channel owner’s age at the time of the event?
    • What are the date(s) of event(s)? we asked for both specific and estimated date(s).
    • How did the attempt/major life event happen? e.g., suicide attempt using medication, or sudden death of a relative.
    The Psychological Assessment step included psychological analysis of predefined relevant factors including suicide risk factors and additional info, and verification of suicide intent. A trained clinical psychologist reviewed the suicide attempt narratives, verifying that the individual describes a suicide attempt rather than self-harm and also conducting a qualitative analysis of risk factor-related information. This annotation was performed only on the Attempted (During) and Attempted (Before) groups and included the following questions:
  • Risk Factors Information:
    • What risk factors for suicide are present (e.g., depression)?
    • What resilience factors are evident (e.g., family support)?
    • What risk group does the individual belong to (e.g., LGBTQ+ individuals)?
    • What appears to be the motivation for posting this video (e.g., raising awareness)?
  • Clinical Information
    • Does the text’s main speaker mention having any mental illness (e.g., PTSD)?
    • Does the text’s main speaker mention being treated by a professional psychologist or psychiatrist and/or being hospitalized?
    • What is the suicide attempt intent (Suicide Attempt, Suicidal Ideation, or Non-Suicidal Self-Harm)?
Figure A.1: The annotation interface from Label-Studio [47], displaying a partial example provided to the annotators. The interface supported two main annotation processes: (1) Non-Expert Assessment, which involved classifying personal channels, estimating demographic information, assessing the relevance of the story, extracting event-related details (e.g., suicide attempt date); and (2) Psychological assessment of suicide attempt narratives (conducted by a clinical psychologist), which included evaluating risk factors, risk groups, mental health illness, and distinguishing suicide attempts from non-suicidal self-harm. Full example of annotation interface is attached in a separate file.
A.4 Matching procedure:
The matching process ultimately identified 95 users as the Control (Matches) group. This group represents a general cohort of the YouTube population matched by different attributes to the Attempted (During) group. While this group was initially collected and matched to the Attempted group, the analysis in the main paper focuses on group-level effects to estimate the prevalence of different topics. The full methodology is illustrated in Figure 5, and the sections below provides both a high-level overview and technical details regarding each step (§A.4).
Overview
Each step in the following section is briefly described, with references to the relevant detailed sections:
(1) We first automatically constructed queries based on treatment group properties (§A.4.1), encompassing channel characteristics such as themes and upload period. This process produced 20–30 candidate matching queries per channel, resulting in a final group of 3,775 channels.
(2) This group was then filtered manually and automatically to identify matched channelsA.4.2) per the inclusion and exclusion criterion detailed in Table A.1, yielding a pool of 597 potential control cases.
We applied a two-step matching algorithm to find an optimal subset of matches:
(3) The first step used propensity score-based matchingA.4.3) [7], which estimates similarity based on numerical and categorical features using a logistic regression model trained on demographic attributes (e.g., age) and channel-related features (e.g., number of uploads). Matching was performed using k-nearest neighbor on the resulting propensity scores.
(4) The second step involved LLM-based refined matchingA.4.4), leveraging GPT-4 [36]. This step incorporated channel content, aiming to control for general themes reflected in the textual content of the channel (e.g., gaming, coding, fashion). Matches were scored on a 1–5 scale based on textual content descriptions, and only channels rated 3 or higher were retained.
This process yielded 95 matched control cases for 42 users who attempted suicide during their upload period, with 1 to 3 matches per treatment channel.
Technical Details
A.4.1 Queries based on treatment group
The initial pool of potential matches was generated by creating YouTube search queries automatically tailored to each user in the Attempted group. These queries incorporated channel-specific features: frequent bi-grams from video titles, video categories, and upload time frames. This process yielded a total of 3,775 channels.
A.4.2 Automatic and manual filtering personal channels
Then, we used the yt-dlp package and downloaded audio files of public videos that were longer than 60 seconds and smaller than 50 MB, and used Whisper (medium) to transcribe the audio. The 3775 channels were later automatically and manually filtered to exclude non-compliant channels and videos (per the inclusion criteria detailed in Table A.1), yielding 597 channels as potential control matches.
A.4.3 Propensity Score-Based Matching:
Propensity Score Estimation:
Propensity score estimation is used to calculate the likelihood of treatment assignment based on observed covariates. A logistic regression model was employed to estimate the propensity score $P(T=t|X)$. Here, $t=1$ represents treatment, Attempted (During), while $t=0$ denotes the Control (Matches) group. The covariates ($X$) consisted of two categories of features:
  • Demographic features: gender and age (manually annotated).
  • Channel-related features: video categories, number of uploads and followers, average upload duration, views, likes, and comments (automatically extracted from YouTube).
Propensity Score Matching:
Based on the estimated propensity scores, k-nearest neighbor matching was applied to identify a set of potential matches for each treatment user. We tested different values of $k$ (3, 5, and 8) and selected $k=8$, which demonstrated higher overall ranks in the subsequent LLM-based evaluation stage.
A.4.4 LLM-based refined matching:
Lastly, we incorporated textual features, channel descriptions, upload titles, and video descriptions, to further refine the matches using GPT-4 [36]. For cost and efficiency, three uploads per channel were preprocessed and summarized using GPT-4. Then, we used GPT-4 again to evaluate and rate each match on a scale of 1 to 5 (using the prompt described in B.1.2), providing explanations for its choices. Matches rated 3 or higher were retained, as they were deemed to be of sufficient quality for analysis. Matches rated below 3 were excluded to reduce noise and improve the reliability of the results.
A.5 Dateset statistics of pre- and post event
The dataset comprised four distinct groups of YouTubers, with varying sample sizes and upload patterns before and after the reference event, described fully in Table A.2. The treatment group, Attempted (During), consists of 42 creators who attempted suicide, of whom 38 had valid upload content after the attempt.
Figure A.2: Box plots representing the number of days between the last upload before the suicide attempt (left) or major life event (right), to the event’s themselves, and afterwards. The y-axis represents logarithmic scale.
In addition, we analyzed the number of days between the last upload before the major life event or suicide attempt, the event itself, and the first upload after the event, in Tables A.3 and A.4. The Attempted (During) group, as seen in Table A.3 has a median time between the last upload before the attempt and the first upload after the attempt of 20 days, while the Control (Major Life Event) group, as seen in Table A.4 had 6 days between the two. To better understand the distribution of inactivity periods, we visualized the gaps using boxplots with a log-scaled y-axis (Figure A.2). While the Attempted group displayed longer gaps overall, we observed that both groups had similarly skewed distributions with multiple long-tail outliers.
We further compared each type of interval between the Attempted (During) and Control (Major Life Event) groups using Welch’s t-test, based on the absolute duration in days and excluding missing values. The analysis revealed that the differences in all three interval types were not statistically significant, indicating similar variability in posting behavior across the two groups.
Group Event Type # Users Pre- Event Post- Event Comments
Mean Median Std Mean Median Std
Attempted (During) Suicide Attempt 42 65 42 80 84 29 148
4 out of 42 users don’t have
valid uploads after the attempt
Attempted (Before) Middle of Uploads 139 102 44 159 98 39 153
midpoint of upload dates was determined
by sorting the unique dates and selecting
the median as the reference point
Control (Major Life Event) Traumatic Event 43 192 88 330 119 50 132
e.g., death or suicide attempt of
parent/child/sibling/close relative
Control (Personal) Middle of Uploads 95 78 42 99 76 42 97
midpoint of upload dates was determined
by sorting the unique dates and selecting
the median as the reference point
Table A.2: Descriptive statistics of YouTubers’ number of uploads pre- and post reference events across four groups - Attempted (During), Attempted (Before), Control (Major Life Events), Control (Personal). The statistics show the mean, median, and standard deviation of uploads in pre- and post-event periods. Comments provide details about event types and reference point calculations.
Interval Description Mean (days) Median (days) SD Min Max
Last upload before $\rightarrow$ First upload after 131.74 59.0 161.48 5 702.0
Last upload before $\rightarrow$ Attempt 83.02 20.5 160.84 1 953.0
Attempt $\rightarrow$ First upload after 72.42 20.5 129.74 1 669
Table A.3: Summary of Temporal Intervals Around Suicide Attempts in the Attempted (During) Group
Interval Description Mean (days) Median (days) SD Min Max
Last upload before $\rightarrow$ First upload after 214.26 39.0 565.95 2 3255
Last upload before $\rightarrow$ Major Life Event 129.77 6.0 506.89 1 3187
Major Life Event $\rightarrow$ First upload after 84.49 17.0 172.63 1 903
Table A.4: Summary of Temporal Intervals Around Traumatic Events for the Control (Major Life Event) Group
Appendix B Appendix B - Procedures
B.1 Topic modeling
Figure B.1: Overview of the Topic Modeling Pipeline. The pipeline includes text segmentation, topic modeling using BERTopic, and video topic distribution to analyze YouTube video content.
B.1.1 Technical Pipeline
The topic modeling pipeline was applied to transcripts from the full-channel videos uploaded by the Attempted (During) and Attempted (Before) groups. Each video was transcribed and preprocessed according to the methods detailed in §A.2.
First, we obtained embeddings of textual segments using Sentence-BERT (SBERT), a modification of the Bidirectional Encoder Representations from Transformers (BERT), network to efficiently create short text embeddings. Uniform Manifold Approximation Projection (UMAP) [31], was used for dimensionality reduction, reducing the embeddings from $300$ to $5$ dimensions. Next, the embeddings were clustered with the Hierarchical Density-based spatial clustering of applications with noise (HDBSCAN) algorithm [30]. HDBSCAN identifies and excludes noisy data points from the clusters, contributing to greater clarity of the topic model. In addition, HDBSCAN provides probability distribution over the topics, as a result of soft clustering (mimicking traditional topic modeling techniques like Latent Dirichlet Allocation (LDA) [5]), in which each document is modeled as a distribution of topics. Finally, each cluster is represented by its top $10$ words and $3$ most representative textual segments, which were used as input to an LLM (specifically LLama-7B [48]), and facilitated in naming the clusters (see B.1.2 for full prompt used in naming the clusters).
While training different topic models, we explored various hyper-parameters and concluded with the following set of parameters:
Table B.1: Model and General Hyper-parameters
Hyperparameter Value
Model meta-llama/Llama-2-7b-chat-hf
Minimum Document Frequency 0.001
Number of Topics auto
N-gram Range (1, 2)
Minimum Topic Size 180
Table B.2: UMAP Hyper-parameters
Hyperparameter Value
Metric cosine
Minimum Distance 0
Low Memory true
Number of Neighbors 15
Number of Components 5
Random State 33
Table B.3: HDBSCAN Hyper-parameters
Hyperparameter Value
Metric euclidean
Minimum Samples 60
Prediction Data true
Minimum Cluster Size 180
Generate Minimum Spanning Tree true
Cluster Selection Method eom
B.1.2 Prompts for Topic Modeling and Matching
This section contains the prompt we used as part of the topic modeling and matching processes. The first prompt used in the LLM-based topic modeling pipeline is described in B.1.2. The LLM (LLama-7B) was provided with the 10 top TF-IDF terms and 3 most representative textual segments, and tasked with generating concise topic labels based on the provided content. The second prompt used in the matching process, as part of the LLM-refined matching by GPT-4 (§A.4.4) and can be seen in B.1.2.
Topic Modeling Naming - <s>[INST] <<SYS>> You are a helpful, respectful and honest assistant for labeling topics. <</SYS>> I have a topic that contains the following documents: - Traditional diets in most cultures were primarily plant-based with a little meat on top, but with the rise of industrial-style meat production and factory farming, meat has become a staple food. - Meat, but especially beef, is the word food in terms of emissions. - Eating meat doesn’t make you a bad person, not eating meat doesn’t make you a good one. The topic is described by the following keywords: ’meat, beef, eat, eating, emissions, steak, food, health, processed, chicken’. Based on the information about the topic above, please create a short label of this topic. Make sure you only return the label and nothing more. [/INST] Environmental impacts of eating meat [INST] I have a topic that contains the following documents: [DOCUMENTS] The topic is described by the following keywords: ’[KEYWORDS]’. Based on the information about the topic above, please create a short label of this topic. Make sure you only return the label and nothing more. [/INST]
Matching Prompt - GPT-4o You will receive information about a treatment YouTuber and <k> potential match YouTubers. For each YouTuber, the following details will be provided: 1. Channel Description: A brief overview of the channel’s content and focus. 2. Summary of Video Titles: A general summary of the types of video titles the channel features. 3. Summary of Video Descriptions: A synopsis of the typical video descriptions found on the channel. 4. Video Tags: Common tags used across the videos. Evaluate the similarity between the treatment YouTuber and each potential match by considering the content, style, and demographic context. Rate each potential match on a scale of 1 to 5, where: 1: Bad match – The channels are very different. 2: Poor match – The channels have some differences. 3: Average match – The channels have some similarities but also notable differences. 4: Good match – The channels are quite similar. 5: Excellent match – The channels are very similar and the individuals seem similar in terms of demographics and context. Provide a brief justification for each rating to explain your reasoning.
B.2 Temporal Alignment
Figure B.2: Overview of the Temporal Analysis Pipeline.
In Figures 9 and 10, we analyzed the temporal dynamics of the topics YouTube Engagement and Mental Health Struggles pre- and post-event. As part of this analysis, we encountered the challenge of aligning different users’ uploads to a mutual timeline. This alignment was performed relative to a specific reference point, which varied by group. For the Attempted (During) and Control (Major Life Event) groups, we used the manually extracted attempt or major life event date, while for the Attempted (Before) and Control (Matches) group, we used the midpoint of upload period. As part of this analysis, we excluded uploads from the attempt/major life event date for all applicable users.
Once reference points were established, we interpolated uploads to weekly values, applying linear interpolation for each topic probability to generate scores for every calendar week and topic. We then standardized the data by up-sampling and down-sampling to obtain 15 samples pre- and post-reference point (total 30). Finally, we aggregated these values across all users, calculating the mean and standard error for each normalized timestamp. This comprehensive approach enabled consistent temporal analysis across diverse user groups, facilitating comparative studies despite varying reference points and upload frequencies.
B.3 Temporal Testing Procedure
In Figures 9 and 10, we performed two types of statistical analyses to compare topic trends:
Procedure 1: Within-Group Comparison (Left-Side Figures)
This procedure compared time steps within the Attempted (During) group to the baseline at $t=0$:
  • Aligned Values: The aligned values of all users in the Attempted (During) group were utilized, with $t=0$ serving as the baseline.
  • Statistical Test: Paired $t$-tests were conducted to compare the aligned values at each time step ($t=1,\ldots,30$) to the baseline values at $t=0$.
  • Post-Hoc Correction: Benjamini-Hochberg (BH) FDR correction was applied to account for multiple comparisons across the 30 time steps.
Procedure 2: Between-Group Comparison (Right-Side Figures)
This procedure compared the Attempted (During) group to each of the three control groups (Attempted (Before), Control (Matches), Control (Major Life Event)) at each time step:
  • Aligned Values: The aligned values of users in the Attempted (During) group were compared to the aligned values of users in the control groups at each time step ($t=1,\ldots,30$).
  • Statistical Test: Welch’s $t$-test, which accounts for unequal variances, was used due to differing sample sizes between groups.
  • Post-Hoc Correction: Benjamini-Hochberg (BH) FDR correction was applied to control for multiple hypothesis testing across 30 time steps and 3 control groups (a total of 90 comparisons).
Both procedures relied on aligned values, as described in §B.2, to ensure comparisons were conducted in a consistent timeline framework.
Appendix C Appendix C - Results
C.1 Topic Modeling of YouTube Channels
Figure C.1: Inter-topic Distance Map showing the results of topic modeling trained on the Attempted (During + Before) group. The figure displays the 100 most frequent topics, with bubble size indicating topic frequency and location determined by the average word embeddings of each topic after dimensionality reduction to 2D using UMAP algorithm. The view is zoomed in to highlight Mental Health topic clusters, with their LLM-generated names visible in blue. The rest of the cluster names were manually assigned for readability.
Our analysis leveraged an LLM-based topic modeling approach to analyze the transcripts from Attempted (Before) and Attempted (During) channels. This analysis resulted in the clustering of these transcripts and the identification of 166 topics, which were automatically named using the Llama-2-7B model [48] (refer to B.1.2 for the full prompt used). The spatial distribution of the topics’ embedding vectors is illustrated in Figure C.1.
Our quantitative analysis of topic quality found that topics with similar names, such as Mental Health Struggles and Mental Health Awareness, were semantically related. This relationship is also reflected in their high cosine similarity [44], with a score of 0.72, indicating that topics with similar LLM-generated names also tend to have similar underlying content.
Additionally, certain topics were user-specific, appearing in less than 10% of the users (e.g., Filipino Culture and Language), while others were more prevalent, appearing in more than 70% of the users’ videos (e.g., YouTube Engagement). A detailed description of the topics, including their names, frequencies, most representative textual segments, and correlations between topics, is provided in the Supplementary Materials Appendix D.
C.2 Topic Modeling of Narrative Suicide Stories
Figure C.2: Attempt Stories Inter-topic Distance Map showing the results of topic modeling trained on the attempt stories. The map visualizes clusters of different topics, named by LLM, where the size of each topic bubble indicates the number of document chunks classified under that topic. The figure also includes zoomed-in views of a cluster of topics centered around mental health experience -(e.g., Mental Health Support, Seeking Help and Support, Religious Mourning and Comfort)
The primary topic model analyzed in this study (§6) was trained on data from the treatment group, including all uploads from both the Attempted (Before) and Attempted (During) groups (including the attempt narrative videos).
As a complementary analysis to the psychological assessment (ref), we trained a separate topic model specifically on the 181 attempt story videos reviewed by the psychologist. We applied a sliding-window text segmentation approach, advancing one sentence at a time, to generate 32,347 overlapping chunks. This method allowed for a more granular representation of the narratives, supporting richer topic detection.
The clustering process yielded 56 distinct topics (Figure C.2), covering themes such as finances, family, personal reflections, friendships, emotional support, crisis management, and resilience. A significant cluster focused on individuals’ mental health journeys, including experiences with self-harm, spirituality, and identity exploration. The clustering results provide a quantitative basis for exploring the interplay between different aspects of individuals’ experiences in the context of suicide attempts.
Future work could explore whether these emergent topics—particularly those not captured in standard clinical assessments—hold predictive value for assessing suicide risk or guiding intervention design.
C.3 Supplementary Statistical Analysis
C.3.1 Bottom-up Approach
Figure C.3 illustrates the word-cloud of the 5 topics identified in this analysis, and Figure C.4 illustrates the average dynamics of the topics Fashion and Style, LGBTQ+ Identity and Acceptance, and Rape and Sexual Assault.
Figure C.3: Word clouds based on the TF-IDF scores of the words in the five topics identified as predictive of the period prior to a suicide attempt (§3.1.1). The size of the words in each cloud is proportional to their TF-IDF score.
Figure C.4 presents the mean aggregated values before and after the reference event, which reveals that the Fashion and Style topic distinguishes the two Attempted groups, showing lower prevalence than the Controls. A review of the demographic information (Table 1) indicates that the proportion of females in the groups does not differentiate between the Attempted and Control groups, suggesting that this factor does not account for the observed difference.
Figure C.4: Fashion and Style (left), LGBTQ+ Identity and Acceptance (middle), and Rape and Sexual Assault (right): Average Comparison Across Groups. The figures present the average topic values across four groups comparing Pre- and Post-Event (suicide attempt, major life event, or synthetic condition). These results highlight the observed trends of lower engagement with Fashion and Style topics in the Attempted groups and higher engagement with LGBTQ+ Identity and Acceptance topics in the Attempted (Before) group. These topics were identified in answering §3.1.1.
C.3.2 Effects Analysis
As part of both the bottom-up and hybrid approaches, we performed a two-way mixed-effects analysis of variance (ANOVA), with time (pre- vs. post-reference point) as a within-subjects factor and group as a between-subjects factor. The four groups included: Attempted (Before), Attempted (During), Control (Major Life Event), and Control (Matches). The goal was to identify topics that differed across groups, changed over time, or showed differential changes across groups, while the main findings are presented in the (§3.1.2) and (§3.4.2).
The results of the two-way mixed-effects ANOVA are summarized in Table C.1, which presents the significant topics identified through the bottom-up (data-driven) approach, and Table C.4, which reflects the outcomes of the hybrid approach that incorporates expert clinical input.
To further investigate these findings, we conducted post-hoc Tukey HSD comparisons for both main and interaction effects:
Group-level effects: Tables C.2 and C.5 report pairwise comparisons between groups, averaged across time. These tests assess whether certain groups consistently exhibited higher or lower topic prevalence, irrespective of temporal dynamics.
Interaction effects: Tables C.3 and C.6 present comparisons of temporal change (i.e., the difference between pre- and post-event scores) across groups. These analyses test whether the direction or magnitude of topic change over time significantly differed between group pairs.
Overall temporal changes: Table C.7 provides the mean topic probabilities for each condition (pre and post) aggregated across all users, offering a descriptive overview of how topic prevalence shifted over time regardless of group membership.
Topics marked with *** were identified by both bottom-up and hybrid approaches.
Table C.1: Two-Way ANOVA summary (Bottom-up-identified topics)
Topic Effect df1 df2 Mean Square F Statistic p-value
Mental Health Struggles *** group 3 311 0.16 4.25 <0.01
condition 1 311 0.11 18.45 <0.01
Interaction 3 311 0.02 3.47 0.02
YouTube Engagement group 3 311 0.59 10.55 <0.01
condition 1 311 0.10 10.18 <0.01
Interaction 3 311 0.05 5.14 <0.01
LGBTQ+ Identity and Acceptance *** group 3 311 0.01 0.28 0.04
Rape and Sexual Assault *** Interaction 3 311 0.001 2.68 0.047
Fashion and Style group 3 311 1.4 27.8 <0.01
Table C.2: Tukey HSD post-hoc Group Comparisons (Bottom-up-identified topics).
Group difference = Group2 mean - Group1 mean.
Topic Group 1 Group 2 Group Difference 95% CI p-adj
Mental Health Struggles *** Attempted (Before) Control (Matches) -0.06 [ -0.11, -0.01 ] 0.01
Attempted (Before) Control (Major Life Event) -0.07 [ -0.13, -0.01 ] 0.03
YouTube Engagement Attempted (Before) Control (Matches) 0.10 [ 0.04, 0.16 ] <0.01
Attempted (Before) Control (Major Life Event) 0.11 [ 0.03, 0.18 ] <0.01
Attempted (During) Control (Matches) 0.12 [ 0.04, 0.2 ] <0.01
Attempted (During) Control (Major Life Event) 0.13 [ 0.03, 0.22 ] <0.01
Fashion and Style Attempted (Before) Control (Matches) 0.18 [0.12, 0.23] <0.01
Attempted (Before) Control (Major Life Event) 0.14 [0.06, 0.21] <0.01
Attempted (During) Control (Matches) 0.16 [0.08, 0.24] <0.01
Attempted (During) Control (Major Life Event) 0.12 [0.03, 0.21] <0.01
Table C.3: Tukey HSD post-hoc Group × Time interaction (Bottom-up-identified topics).
Group and Time Difference = (Group2 post - pre) - (Group1 post - pre).
Topic Group 1 Group 2 Group and Time Difference 95% CI p-adj
Mental Health Struggles *** Attempted (During) Attempted (Before) -0.06 [-0.12, -0.08] <0.01
YouTube Engagement Attempted (During) Attempted (Before) -0.08 [-0.14, -0.01] 0.02
Attempted (During) Control (Matches) -0.11 [-0.18, -0.04] <0.01
Attempted (During) Control (Major Life Event) -0.08 [-0.17, -0.01] 0.04
Table C.4: Two-Way ANOVA summary (Hybrid-identified topics)
Topic Effect df1 df2 Mean Square F Statistic p-value
Autism and Disability condition 1 311 0.004 6.23 0.01
Underage Smoking and Drug Use condition 1 311 0.006 13.63 <0.001
Homelessness condition 1 311 0.040 17.87 <0.001
Interaction 3 311 0.009 3.95 0.009
Anxiety and Panic Attacks group 3 311 0.008 4.28 0.006
Interaction 3 311 0.002 3.45 0.017
Social Media Addiction group 3 311 0.008 3.72 0.012
condition 1 311 0.011 8.86 0.003
Addiction Recovery condition 1 311 0.030 19.54 <0.001
Sleep Disorders and Habits group 3 311 0.007 3.42 0.018
condition 1 311 0.011 10.12 0.002
Trauma-Induced Mental Health Disorders condition 1 311 0.009 9.43 0.002
Medication Use and Addiction group 3 311 0.016 3.66 0.013
condition 1 311 0.007 10.47 0.001
Table C.5: Hybrid Approach: Tukey HSD Multiple Comparison of Groups
Topic Group 1 Group 2 Group Difference 95% CI p-adj
Anxiety and Panic Attacks Attempted (Before) Control (Matches) -0.01 [ -0.02, -0.003 ] 0.01
Medication Use and Addiction Attempted (Before) Control (Matches) -0.02 [ -0.03, -0.002 ] 0.03
Table C.6: Tukey HSD post-hoc Group Comparisons (Hybrid-identified topics).
Group Difference = Group2 mean - Group1 mean.
Topic Group 1 Group 2 Group and Time Difference 95% CI p-adj
Homelessness Attempted (Before) Control (Matches) 0.02 [0.001, 0.05] 0.04
Attempted (Before) Control (Major Life Event) 0.03 [0.004, 0.06] 0.02
Anxiety and Panic Attacks Attempted (Before) Control (Major Life Event) 0.02 [0.01, 0.04] 0.01
Table C.7: Mean topic probabilities over all users pre- and post-event (Bottom-up and Hybrid)
Topic Mean (Pre) Mean (Post)
Medication Use and Addiction 0.03 0.03
Homelessness 0.05 0.07
Trauma-Induced Mental Health Disorders 0.03 0.04
Fashion and Style 0.16 0.19
Addiction Recovery 0.05 0.07
Underage Smoking and Drug Use 0.01 0.02
Sleep Disorders and Habits 0.04 0.05
YouTube Engagement 0.41 0.44
Autism and Disability 0.03 0.03
Mental Health Struggles 0.17 0.20
C.3.3 Control Measures Analysis
Figure C.5: Temporal trends around the start of the COVID-19 pandemic in the USA (March 1, 2020), and around Reference Event (suicide attempt, major life event and midpoint of upload) in the topics COVID-19 Pandemic (top) and Mental Health Struggles (bottom).
To corroborate that the observed topic shift was specifically associated with suicide attempts and not driven by other external factors, we examined topic dynamics around the COVID-19 pandemic, which served as an absolute reference point unrelated to individual users. We performed the same between analyses as in 3.2.
This analysis revealed an expected pattern as seen in Figure C.5: both the treatment and control groups exhibited an sudden increase in the topic COVID-19 Pandemic around the pandemic’s onset (March 1st 2020), that didn’t show any significant difference between the groups, and a gradual increase in the topic Mental Health Struggles, that is different in the temporal dynamics than the pattern detected by the suicide attempt.
Appendix D Appendix D - Topic Information
This section includes two tables with prevalent topics identified in this study. Table LABEL:tab:topic_info describes the topics identified in our bottom-up (§3.1) and hybrid approaches (§3.4). Table D.2 presents topics that are highly correlated and have similar semantic meaning.
Table D.1: Overview of topic information, including: Count - the number of transcript segments assigned to each topic, Llama2 Name - the name generated by the Llama-7b model (see Section B.1.2), Representative Transcript Segment - the transcript segment closest to the topic’s center and most representative of it, and Top Words - the 10 words with the highest TF-IDF scores. Red topic names indicate topics that were identified in our bottom-up approach. Blue topics were found in our hybrid approach, and brown indicates topics that were found by both approaches. Corresponding * indicates significance with regards to temporal changes in Attempted (During).
Count Representative Transcript Segment Llama2 Name Top Words
11815 quite what to do and I turn to self harm and suicide attempts to cope with mine which was not the best way to cope but I didn’t know what else to do, I was hopeless and I just felt like so horrible and alone and I never want to go back to feeling like that. You need to think that number one it’s not the end of the world, people do get back together, Mental Health Struggles * suicide,self harm,harm, depression, suicidal,bipolar, trauma,self, mental, scars
19822 again y’all i enjoyed doing this video for y’all hope that you enjoyed watching it if you made it to the end give me those thumbs up make sure you like comment subscribe to my channel hit the bell twice to be notified if you’re an old subscriber make sure you’re on share my videos and like i always always say god only gives you one life y’all please live it bye y’all YouTube Engagement * youtube,subscribers, thumbs, channel,subscribe,videos, hey people, subscribe channel,views,live
19387 As you can see, like this white part, it’s a little bit like, like just like the strings you can like see through it kind of. So, that’s pretty awesome. I absolutely love this bikini. And honestly, I didn’t think this color would look good on me, but I actually really, really like it. Oh, and by the way, these bags are cute. Like they’re very good quality bags and like they look really nice. Fashion and Style wear,size,dress, jeans,cute,shirt, wearing,really nice,shorts, pants
6729 I want to try to come up against the scope. In Jesus’ name, Lord, I pray and Lord, I thank you that you get the glory from this scope, God. You get the glory from this Lord, not me. In Jesus’ name, I pray. Amen. Okay and ignore my hair. I just got done washing my hair because my, my dandruff was acting up so don’t mind me but as I was, you know, looking Religious Prayer and Faith jesus,god, lord, church, pray, faith, prayer, praying, jesus lord, spirit, god god, bible, religion,spiritual,verse
10944 What’s up guys? What is life as a cocaine addict like? What is it like to be a cocaine addict? I was addicted to cocaine. My addiction was to the point where it was one to two eight balls of coke a day. If you don’t know what an eight ball of coke is, that is 3.5 to almost four grams of coke, depending on if you know the dealer or not and depending on what it’s cut with. Addiction Recovery sober,nofap, addiction, cocaine, drinking, addict, alcohol,sobriety, relapse,coke
1795 uh feeling tired like I would go to bed early and I just wanted to sleep like I really really needed to sleep week nine day three I was so happy because I’ve been waking up really late every day like almost 9 a.m because I feel so tired and so fatigued I made my bed this morning for the first time since like a week because I I didn’t even have energy to make the Sleep Disorders and Habits sleep,wake, waking, bed, morning, sleeping, morning routine, hours sleep, sleep schedule, hours
3363 Just look at it, man. Right now what we have going on on the internet, we got black men against black women. Then on the streets we got black men against black men, and then black women against black women. We all talk about wanting unity, but where is it? We definitely have to do better. Black community black,black women,black men,black people,black woman,white,racist,black man,white people,women
4124 If you like this video, I would love it if you’d give it a thumbs up. And if you’re interested in all things autism, I’d love it if you’d hit subscribe and I’ll see you tomorrow. Bye. Autism and Disability autism, autistic, trans, transgender, disabled, things autism,interested things, autism love, thumbs interested, love hit
952 A lot of people just say that I’m gay or they assume I’m gay when I’m bi. They’re like oh no you can’t be bi you’re so gay. I’m bi I think. Seven what’s your favorite thing about the LGBT community? I just love how we’re all really caring for each other. There are hateful people but most of the LGBT community is really loving and accepting LGBTQ+ Identity and Acceptance * gay,bisexual,lesbian,lgbtq, sexuality,gay people,homophobic,queer, pansexual,lgbtq community,lgbt, straight,gay gay,bi,gays
221 Okay, yes, I know that there are a lot of women out here who are falsely accusing men of sexual assault and rape. I know this, I know this, I know this, but don’t go into the situation thinking that’s what’s going on. You have to believe a female when she says that she’s been raped. Rape and Sexual Assault * rape,raped , consent,considered rape,rape know, consented, victims,rapist,assault,sexual assault, men raped,rape
1041 I’m going to pop in something really sour into my mouth, or I’m going to smell something that’s going to get my brain out of here. Any tips on, uh, anxiety panic attacks, panic attacks are hard to ground yourself from and they’re hard to stop yourself. There is huge differences between anxiety attacks and panic attacks. One of the biggest ones is there’s no trigger necessarily behind a panic attack. Anxiety and Panic Attacks * anxiety,panic,attacks,panic attacks,panic attack,attack,anxiety attack,anxious,anxiety attacks,anxiety like
763 and year 9, not a lot of people knew about me. I was just kind of this loser. I was called emo, but no one really bullied me and even if people tried to, it never fazed me. I’ve never been bullied in that sense. When I was in school I’ve never been bullied and it doesn’t mean that other people haven’t tried to bully me, it’s just that I don’t take it as bullying. Bullying bullied, bullying, bully, bullies, like bullied, bullied like, getting bullied, ve bullied, school, cyber
194 Just tell them what you just told us. Okay. This is a situation that most people don’t see when they see the homeless people. They see us as a problem. You have homeless people that are on the street as well as in the shelters, but the shelters don’t usually work out for us, so we end up coming back out to the street. If you get in an argument with somebody, you’re kicked out of the shelter, Homelessness homeless,homeless people,homeless person,people homeless,say homeless,homeless like,homeless population,tourists,know homeless
1262 finding out what works for you when it comes to the time of medication because you can hear horror stories of other people with Adderall I mean I know people that have said that they popped Adderall like Skittles and I don’t comprehend that because I take 10 milligrams in the morning I take 10 milligrams in the afternoon and there’s other people out there that are like you Medication Use and Addiction adderall, xanax, medication, effects, prescribed, milligrams, 10 milligrams, adhd, ritalin, prozac
250 there’s signs out front that say, you know, a vet lives here, please be cautious and courteous with your fireworks. For some, it’s slamming doors, it’s loud voices, it’s slamming objects. That triggers your flashback. That sparks and flares up their PTSD. There is that big difference. So let me really break down the differences between the two. PTSD, one or few trauma, short-lived trauma. Visual and somatic flashbacks, avoidance of triggers, isolating, nightmares about trauma, Trauma-Induced Mental Health Disorders ptsd, flashbacks, cptsd, triggers, smell blood, emdr therapy, emdr, complex ptsd, ptsd ptsd
192 Honest chat. Honest chat. So one of you actually asked the question, what is BPD and is there a way to get over it? So I thought that would be a good starting point to explain what BPD is. BPD is what I’ve been diagnosed with along with PTSD, anxiety and depression. BPD is borderline personality disorder. It can also be called EUPD, emotionally unstable personality disorder. Personally I think that’s a little bit offensive. Living with Borderline Personality Disorder (BPD) bpd, people bpd, personality disorder, borderline personality, borderline, personality, disorder, bpd like, diagnosed bpd, diagnosed
3363 So Amanda, how are you today? I’m doing okay. So I can tell you a little bit about Obsessive-Compulsive Disorder, or OCD. It’s a mental illness and it’s one of the anxiety disorders. And so there’s two parts. There’s the obsessions and there’s the compulsions. So the obsessions, those are repetitive and intrusive thoughts. It can be urges or images that cause a lot of anxiety in an individual. And then the second part are the compulsions. Obsessive-Compulsive Disorder (OCD) ocd, compulsions, intrusive, intrusive thoughts, people ocd, ocd like, obsessions, ocd just, obsessive compulsive, obsessive
1036 Hi guys welcome to educating Shani I’m Shani and I’m recovering from an eating disorder. Hi I’m Michelle. Wait what am I supposed to say? Say hi Shani. Hi Shani. Hi! Welcome to… Start over. Hi guys welcome to educating Shani I’m Shani and I’m recovering from an eating disorder. I’ve been in this chair for like two. Yeah well it won’t be two hours after editing. Eating Disorder Recovery shani,shanny,hi shani,welcome educating,educating,disorder hi,educating shani,shani shani,shani hi,shani recovering
219 She becomes a rebellious sure goes off to college and Cal basically just goes to high school and does his own thing Amore stays in it or Sonny stays in the crib and that’s it right Basil It’s so weird cuz Basil I feel like Basil Basil and Amore saw Mary kill herself or like walked in on Mary Which is why it’s like they’re both kind of like I would say traumatized a little bit Loss and Trauma in a Small Community basil,aubrey, mary, sonny, amore, kel, mary died, photo album, sunny, hero, basil basil, mori,right aubrey, aubrey aubrey, hector
186 I make an effort to never do anything like that. I do not want to set unrealistic standards for you guys. And when I say unrealistic, I don’t mean that those things aren’t attainable. What I mean is that you shouldn’t be aiming for that. You know, you shouldn’t be striving for perfection because perfection doesn’t exist. Your imperfections are what make you perfect. Perfectionism and the Pursuit of Imperfection perfect, perfection, imperfect, imperfections, perfectionism, perfect perfect, perfectionist, know perfect, perfect person, perfect don
Table D.2: Correlation Matrix of 30 Topics Based on Cosine Similarity: The matrix, computed using cosine similarities of the topic embeddings, demonstrates that topics with high correlation are also semantically similar, as anticipated.
Topic 1 Topic 2 Correlation
Online Therapy Mental Health Struggles 0.756
Mental Health Struggles Online Therapy 0.756
Singing and Going Mental Health Struggles 0.750
Mental Health Struggles Singing and Going 0.750
YouTube Engagement Vlog End and Daily Uploads 0.744
Vlog End and Daily Uploads YouTube Engagement 0.744
Underage Smoking and Drug Use Addiction Recovery 0.743
Addiction Recovery Underage Smoking and Drug Use 0.743
Social Media Addiction Mindful Social Media Use 0.742
Mindful Social Media Use Social Media Addiction 0.742
Tumultuous online interaction Loss and Trauma in a Small Community 0.741
Loss and Trauma in a Small Community Tumultuous online interaction 0.741
Mental Health Awareness YouTube Engagement 0.728
YouTube Engagement Mental Health Awareness 0.728
Anxiety and Panic Attacks Fears and Growth 0.725
Fears and Growth Anxiety and Panic Attacks 0.725
Mental Health Awareness Mental Health Struggles 0.723
Mental Health Struggles Mental Health Awareness 0.723
Self-Confidence and Overcoming Adversity Self-perception and beauty standards 0.718
Self-perception and beauty standards Self-Confidence and Overcoming Adversity 0.718
YouTube Engagement Social media content creation and scheduling 0.716
Social media content creation and scheduling YouTube Engagement 0.716
Yoga Practice and Philosophy Mindfulness and Meditation 0.715
Mindfulness and Meditation Yoga Practice and Philosophy 0.715
Christmas Tree Decorating Toxic Christmas 0.711
Toxic Christmas Christmas Tree Decorating 0.711
Vaccine Beliefs and Misconceptions COVID-19 Pandemic 0.709
COVID-19 Pandemic Vaccine Beliefs and Misconceptions 0.709
Mental Health Struggles Emotional Resistance 0.709
Emotional Resistance Mental Health Struggles 0.709